# Modeling Clinical Uncertainty in Radiology Reports: from Explicit Uncertainty Markers to Implicit Reasoning Pathways

### v1

@@ 1.  Motivation and Related Work
Radiology reports play a central role in clinical decision-making, serving as the primary medium through which radiologists communicate their interpretations and diagnostic impressions to referring physicians. These reports influence downstream diagnostic reasoning, and often determine treatment trajectories. As AI models have been increasingly used for both automated radiology report interpretation and generation, structuring frameworks have been developed to convert free-text reports into machine-readable formats that can be used for training and evaluation (lunguage; chest_imagenome; radgraph; radgraph2). A key challenge that needs to be addressed is that radiology reports inherently contain uncertainty. This uncertainty arises in two distinct forms: explicit uncertainty, expressed directly through the language radiologists use to qualify their findings or diagnoses (communicating_uncertainty; communication_doubt_radiology), and implicit uncertainty, which emerges from the selective and often incomplete nature of what is recorded in the report (linguistic_uncertainty_clinical). Figure 1 shows how these two forms of uncertainty appear in chest X-ray (CXR) reports, which are especially prone to contain uncertainty (uncertainty_nlp; chexpert).
@@
Explicit uncertainty arises when radiologists convey doubt about the presence or absence of a finding or diagnosis, typically through hedging phrases such as “probably”, “possible”, “suggesting”, or “may represent”, among many others. Prior work examining these expressions has highlighted their prevalence, revealing the mention of uncertain diagnoses in a high proportion of CXR reports (chexpert; lunguage). The use of hedging in radiology reports is deliberate and meaningful: it enables radiologists to communicate diagnostic uncertainty that is crucial for appropriate clinical interpretation and decision-making (quantitative_uncertainty; communicating_uncertainty). Automated systems should similarly be designed to recognize and account for this uncertainty, just as human readers of radiology reports are trained to do.
@@
Previous work has extracted uncertainty using rule-based systems (chexpert; mimic_cxr_jpg; cad_chest) in ternary classification settings (positive / uncertain / negative). These approaches are limited in scope, using predefined vocabularies of hedging phrases that trigger an uncertain label when detected in the report. In this work, we go beyond discrete labeling by quantifying uncertainty as a continuous probability between 0 and 1, taking into account the specific hedging phrases and their sentence-level context (see Figure 1). Earlier attempts to assign probabilistic meanings to hedging expressions have typically relied on human judgments, asking experts to rate phrases or position them along a scale of certainty (communication_doubt_radiology; diagnostic_uncertainty_radiology). However, these approaches proved unreliable, as the interpretation of hedging varies widely across radiologists. To address this, we introduce an automated approach that estimates the probability of a finding by leveraging large language models to perform pairwise comparisons of uncertainty expressions, constructing a relative ranking that is then mapped to a continuous probability.
@@
Beyond overt language, reports also embody a subtler form of implicit uncertainty that has received little systematic attention. Radiologists frequently omit portions of their diagnostic reasoning, documenting only key findings or final impressions to maintain conciseness and ensure that the main message is easily understood by the reader (implicit_cognitive). For example, a report may state “congestive heart failure” without mentioning sub-findings that are commonly associated with the condition, such as “consolidation” or “cardiomegaly”. Consequently, it is often unclear whether unmentioned findings are truly absent or simply unrecorded, even though knowing this distinction is crucial, since misinterpreting unrecorded evidence as true absence can systematically bias the data and distort the inferred diagnostic reasoning (implicit1; implicit_cognitive; implicit_bias; implicit_pitfall).
@@
In other words, implicit uncertainty does not stem from linguistic ambiguity, but from the selective and incomplete nature of clinical reporting (linguistic_uncertainty_clinical) – what evidence is stated, abstracted, or left unstated within the report. Disentangling these possibilities requires contextual understanding of diagnostic logic and domain expertise, making implicit uncertainty difficult to model and largely unaddressed in prior research. Understanding and modeling these uncertainties is not only critical for structuring radiology reports but also for developing AI systems that faithfully capture the reasoning process of radiologists. By explicitly representing both what is uncertain and what is implied, such structured resources enable more reliable training and evaluation of medical AI models in uncertainty-aware report generation and interpretation.
@@
  • Releasing Lunguage++, which extends the Lunguage dataset of structured radiology reports (lunguage) by incorporating our techniques for capturing explicit and implicit uncertainty.
@@ 2.  Lunguage Dataset
We demonstrate our methods on the Lunguage dataset (lunguage), a benchmark dataset containing 1,473 annotated CXR reports from the MIMIC-CXR dataset (mimic_cxr). Each report has been structured into fine-grained (finding, relation, attribute) triplets, where findings represent core clinical concepts (e.g. “opacity”, “pneumonia”), and relations specify their contextual links (e.g., “location”, “severity”) with corresponding attributes (e.g., “right lower lobe”, “moderate”). We use the Lunguage dataset because of its extensive granularity in the included findings, relations and attribute types. Furthermore, each annotated finding includes a binary label (tentative and definitive) quantifying the confidence expressed by the radiologist – hedging phrases (e.g., “suggests”, “cannot exclude”) lead to a tentative label, while findings that lack such phrases are labeled definitive.
@@ 3.  Explicit uncertainty
Previous attempts to map hedging phrases to probabilities relied on expert ratings of individual phrases, an approach shown to be unreliable due to inconsistency across experts (communication_doubt_radiology; diagnostic_uncertainty_radiology). Furthermore, the context of the sentence beyond individual phrases should be taken into account: “probably pneumonia” conveys a different certainty than “probably pneumonia given patient history”, even though both use the phrase “probably”. To address these limitations, we adopt a different strategy, which relies on large language models (LLMs) to perform in-context pairwise comparisons between sentences conveying uncertainty. An overview of our framework is shown in Figure 2. We now describe each step of the process in detail. Additional details can be found in Appendix A.
@@ 3.1.  Extracting a vocabulary of common hedging phrases
For each finding-sentence pair, we automatically extract the hedging phrases that convey uncertainty about the finding $f_{i}$. We do this by prompting Gemini (gemini). The main part of the prompt is shown in Listing 1, while the full prompt contains a system message, ten in-context examples and additional instructions (see Appendix A.1). The prompt specifies the finding $f_{i}$ to avoid extraction of hedging phrases from the sentence which have nothing to do with that particular finding. We include in our vocabulary all hedging phrases that were extracted ten times or more, resulting in a vocabulary of 42 hedging phrases. Each phrase is associated with a list of finding-sentence pairs where it was extracted. The five most common extracted hedging phrases include or (373 times), likely (239 times), may (215 times), suggesting (74 times), and cannot be excluded (71 times); the full vocabulary is found in Appendix A.1.
@@ 3.2.  Building a reference ranking of hedging phrases
We draw inspiration from ranking systems in competitive gaming, where player skill is inferred from the outcomes of matches. Specifically, we employ the TrueSkill algorithm (trueskill), a Bayesian rating system that updates each item’s mean skill level $\mu$ based on pairwise comparisons with other items. This way, TrueSkill efficiently infers a global ranking that reflects the relative skill level $\mu$ of each item. In our case, the items are hedging phrases.
@@
To obtain a reliable and well-calibrated ranking, a large number of pairwise comparisons between phrases is required. We therefore leverage LLMs to perform these comparisons automatically, employing three general-purpose LLMs (Gemini (gemini), GPT-4o (gpt_4o), and Claude (claude)) and one domain-specific medical LLM (MedGemma (medgemma)). [^2] To deal with the sensitive nature of the reports in Lunguage we (i) ran a HIPAA-compliant GPT-4o model provided by Azure, (ii) revoked data retention rights for Gemini and Claude, and (iii) ran MedGemma locally. Since the preferred phrase can vary depending on the sentence context, we conduct ten comparisons for each pair of the 42 hedging phrases, where in each comparison, we randomly sample a sentence in which the phrase occurs. We ask each LLM to identify which phrase conveys greater certainty about the presence of a finding, ensuring that such phrases will eventually receive a higher $\mu$. The main portion of the prompt is shown in Listing 2; the complete prompt additionally includes a system instruction, detailed task description, and three in-context examples (Appendix A.2). To ensure a more neutral comparison, the referenced finding $f_{i}$ within each sentence is masked as <finding>.
@@
With the reference ranking established, we fit each of the 2,066 tentative finding-sentence pairs from the Lunguage dataset into it. Note that we cannot simply use the hedging phrases extracted from the sentence to assign a rank directly, because (i) some phrases are too rare to appear in the reference vocabulary, and (ii) the context in which a phrase occurs can significantly alter its implied certainty. Here, we once again draw inspiration from competitive gaming: when a new player enters a game, TrueSkill quickly estimates their rank by identifying existing players with similar skill levels $\mu$. This is done by selecting opponents with the highest draw probability, iteratively playing those games, and updating the skill level of the new player based on the outcome of each game (trueskill).
@@ 4.  Implicit uncertainty
Radiology reports often encode diagnostic reasoning implicitly (implicit_bias; implicit_cognitive; implicit_pitfall; implicit1), documenting high-level diagnoses while omitting intermediate supporting findings. Although such abstraction enhances efficiency, it introduces ambiguity arising from unmentioned findings. To resolve this, we propose a Pathway Expansion Framework that reconstructs omitted diagnostic evidence by systematically expanding structured findings along predefined diagnostic pathways. This framework yields Lunguage++, an expanded version of Lunguage, integrating explicit and implicit reasoning into a unified structured representation (Figure 4).
@@ 4.1.  Diagnostic Pathway Construction
We formalize radiologists’ implicit diagnostic reasoning into explicit, machine-friendly formats that describe how each diagnosis decomposes into its characteristic radiographic findings. To preserve interpretability and clinical fidelity, we established three principles through expert consensus. (i) Exclusive mutual independence ensures that each pathway differs from others by at least one defining piece of evidence. (ii) Specificity ensures that observations directly contributing to diagnostic differentiation are included, emphasizing findings that are concrete and fine-grained. (iii) High-certainty features retain only clinically consistent and reliable findings to minimize ambiguity and preserve clarity. Pathways were then constructed through a two-stage expert-in-the-loop process grounded in these principles. A radiologist and an AI researcher first defined subfindings for 14 common CXR diagnoses, referencing established radiological interpretation principles (goodman2014felson; webb2011thoracic) to identify diagnostic criteria, imaging patterns, and hierarchical relations among findings. These preliminary pathways were subsequently reviewed and refined through consensus by another radiologist and an oncologist.
Pathways are represented as Directed Acyclic Graphs (DAGs) (pearl2009causality), capturing how higher-level diagnoses require lower-level manifestations within hierarchical reasoning. An example of a DAG corresponding to (part of) the diagnostic pathways for congestive heart failure (CHF), pulmonary edema, pleural effusion, and cardiomegaly is shown in Figure 4. Each node corresponds to an observable finding, and directed edges $(u,v)$ represent diagnostic relations $p_{path}(v|u)$ established through radiologist consensus, indicating that the sub-finding $v$ is typically observed when its parent $u$ is present, with likelihood greater than 0.8. This top-down structure models the cascading sequence of radiographic evidence that underlies diagnostic inference. Distinct pathways are defined for each diagnosis and further refined by imaging view and patient position, since these factors influence the findings one expects to observe. Table 3 reports statistical summaries of the diagnostic pathways, while the full set of pathways is provided in Appendix B.1 and on our Github repository.
@@ 4.2.  Pathway Expansion Framework
Building upon the diagnostic pathways, we design a Pathway Expansion Framework that implements the top-down reasoning process of radiological interpretation to fill in missing sub-findings. Given structured findings from Lunguage (lunguage), the framework reconstructs omitted intermediate findings by expanding each diagnosis along its predefined diagnostic pathway, following the recursive procedure detailed in Algorithm 1. It consists of four stages, each of which ensure that reconstructed structures remain clinically faithful and logically consistent.
@@ 1. Finding Deduplication
To ensure that subsequent reasoning operates on a consistent and non-redundant set of findings, overlapping or synonymous mentions referring to the same diagnostic concept at the same location (e.g., “opacity in the basal segment of the lung” and “lung base opacity”) are identified and merged within each report. Pairwise cosine similarity is computed between clinical embeddings (BioLORD (remy2023biolord)) of all report findings, where each finding is linearized into a phrase combining its “entity, location, attributes” following the LunguageScore (lunguage) formulation. Pairs exceeding a similarity threshold of 0.9 are merged into a single finding, while manually defined blacklist pairs (e.g., left vs. right) are excluded to prevent erroneous merges of semantically distinct findings. This precision-oriented threshold and manual blacklist correction minimize erroneous merges while ensuring that each finding is represented only once before expansion. The detailed blacklist pairs are provided in Appendix B.2.1.

### v2

# arXiv:2511.04506v2[cs.CL] 27 Feb 2026

@@ 1.  Motivation and Related Work
Radiology reports play a central role in clinical decision-making, serving as the primary medium through which radiologists communicate their interpretations and diagnostic impressions to referring physicians. These reports influence downstream diagnostic reasoning, and often determine treatment trajectories. As AI models have been increasingly used for both automated radiology report interpretation and generation, structuring frameworks have been developed to convert free-text reports into machine-readable formats that can be used for training and evaluation (Moon et al., 2026; Wu et al., 2021; Jain et al., 2021; Khanna et al., 2023). A key challenge that needs to be addressed is that radiology reports inherently contain uncertainty. This uncertainty arises in two distinct forms: explicit uncertainty, expressed directly through the language radiologists use to qualify their findings or diagnoses (Bruno et al., 2017; Hobby et al., 2000), and implicit uncertainty, which emerges from the selective and often incomplete nature of what is recorded in the report (Turner et al., 2021). Figure 1 shows how these two forms of uncertainty appear in chest X-ray (CXR) reports, which are especially prone to contain uncertainty (Callen et al., 2020; Irvin et al., 2019).
@@
Explicit uncertainty arises when radiologists convey doubt about the presence or absence of a finding or diagnosis, typically through hedging phrases such as “probably”, “possible”, “suggesting”, or “may represent”, among many others. Prior work examining these expressions has highlighted their prevalence, revealing the mention of uncertain diagnoses in a high proportion of CXR reports (Irvin et al., 2019; Moon et al., 2026). The use of hedging in radiology reports is deliberate and meaningful: it enables radiologists to communicate diagnostic uncertainty that is crucial for appropriate clinical interpretation and decision-making (Reiner, 2018; Bruno et al., 2017). Automated systems should similarly be designed to recognize and account for this uncertainty, just as human readers of radiology reports are trained to do.
@@
Previous work has extracted uncertainty using rule-based systems (Irvin et al., 2019; Johnson et al., 2019b; Zhang et al., 2023) in ternary classification settings (positive / uncertain / negative). These approaches are limited in scope, using predefined vocabularies of hedging phrases that trigger an uncertain label when detected in the report. In this work, we go beyond discrete labeling by quantifying uncertainty as a continuous probability between 0 and 1, taking into account the specific hedging phrases and their sentence-level context (see Figure 1). Earlier attempts to assign probabilistic meanings to hedging expressions have typically relied on human judgments, asking experts to rate phrases or position them along a scale of certainty (Hobby et al., 2000; Shinagare et al., 2019). However, these approaches proved unreliable, as the interpretation of hedging varies widely across radiologists. To address this, we introduce an automated approach that estimates the probability of a finding by leveraging large language models to perform pairwise comparisons of uncertainty expressions, constructing a relative ranking that is then mapped to a continuous probability.
@@
Beyond overt language, reports also embody a subtler form of implicit uncertainty that has received little systematic attention. Radiologists frequently omit portions of their diagnostic reasoning, documenting only key findings or final impressions to maintain conciseness and ensure that the main message is easily understood by the reader (Lee et al., 2013). For example, a report may state “congestive heart failure” without mentioning sub-findings that are commonly associated with the condition, such as “consolidation” or “cardiomegaly”. Consequently, it is often unclear whether unmentioned findings are truly absent or simply unrecorded, even though knowing this distinction is crucial, since misinterpreting unrecorded evidence as true absence can systematically bias the data and distort the inferred diagnostic reasoning (Gärtner et al., 2020; Lee et al., 2013; Gunderman, 2009; Berlin, 2000).
@@
In other words, implicit uncertainty does not stem from linguistic ambiguity, but from the selective and incomplete nature of clinical reporting (Turner et al., 2021) – what evidence is stated, abstracted, or left unstated within the report. Disentangling these possibilities requires contextual understanding of diagnostic logic and domain expertise, making implicit uncertainty difficult to model and largely unaddressed in prior research. Understanding and modeling these uncertainties is not only critical for structuring radiology reports but also for developing AI systems that faithfully capture the reasoning process of radiologists. By explicitly representing both what is uncertain and what is implied, such structured resources enable more reliable training and evaluation of medical AI models in uncertainty-aware report generation and interpretation.
@@
  • Releasing Lunguage++, which extends the Lunguage dataset of structured radiology reports (Moon et al., 2026) by incorporating our techniques for capturing explicit and implicit uncertainty.
@@ 2.  Lunguage Dataset
We demonstrate our methods on the Lunguage dataset (Moon et al., 2026), a benchmark dataset containing 1,473 annotated CXR reports from the MIMIC-CXR dataset (Johnson et al., 2019a). Each report has been structured into fine-grained (finding, relation, attribute) triplets, where findings represent core clinical concepts (e.g. “opacity”, “pneumonia”), and relations specify their contextual links (e.g., “location”, “severity”) with corresponding attributes (e.g., “right lower lobe”, “moderate”). We use the Lunguage dataset because of its extensive granularity in the included findings, relations and attribute types. Furthermore, each annotated finding includes a binary label (tentative and definitive) quantifying the confidence expressed by the radiologist – hedging phrases (e.g., “suggests”, “cannot exclude”) lead to a tentative label, while findings that lack such phrases are labeled definitive.
@@ 3.  Explicit uncertainty
Previous attempts to map hedging phrases to probabilities relied on expert ratings of individual phrases, an approach shown to be unreliable due to inconsistency across experts (Hobby et al., 2000; Shinagare et al., 2019). Furthermore, the context of the sentence beyond individual phrases should be taken into account: “probably pneumonia” conveys a different certainty than “probably pneumonia given patient history”, even though both use the phrase “probably”. To address these limitations, we adopt a different strategy, which relies on large language models (LLMs) to perform in-context pairwise comparisons between sentences conveying uncertainty. An overview of our framework is shown in Figure 2. We now describe each step of the process in detail. Additional details can be found in Appendix A.
@@ 3.1.  Extracting a vocabulary of common hedging phrases
For each finding-sentence pair, we automatically extract the hedging phrases that convey uncertainty about the finding $f_{i}$. We do this by prompting Gemini (Comanici et al., 2025). The main part of the prompt is shown in Listing 1, while the full prompt contains a system message, ten in-context examples and additional instructions (see Appendix A.1). The prompt specifies the finding $f_{i}$ to avoid extraction of hedging phrases from the sentence which have nothing to do with that particular finding. We include in our vocabulary all hedging phrases that were extracted ten times or more, resulting in a vocabulary of 42 hedging phrases. Each phrase is associated with a list of finding-sentence pairs where it was extracted. The five most common extracted hedging phrases include or (373 times), likely (239 times), may (215 times), suggesting (74 times), and cannot be excluded (71 times); the full vocabulary is found in Appendix A.1.
@@ 3.2.  Building a reference ranking of hedging phrases
We draw inspiration from ranking systems in competitive gaming, where player skill is inferred from the outcomes of matches. Specifically, we employ the TrueSkill algorithm (Herbrich et al., 2006), a Bayesian rating system that updates each item’s mean skill level $\mu$ based on pairwise comparisons with other items. This way, TrueSkill efficiently infers a global ranking that reflects the relative skill level $\mu$ of each item. In our case, the items are hedging phrases.
@@
To obtain a reliable and well-calibrated ranking, a large number of pairwise comparisons between phrases is required. We therefore leverage LLMs to perform these comparisons automatically, employing three general-purpose LLMs (Gemini (Comanici et al., 2025), GPT-4o (Hurst et al., 2024), and Claude (Anthropic, 2025)) and one domain-specific medical LLM (MedGemma (Sellergren et al., 2025)). [^2] To deal with the sensitive nature of the reports in Lunguage we (i) ran a HIPAA-compliant GPT-4o model provided by Azure, (ii) revoked data retention rights for Gemini and Claude, and (iii) ran MedGemma locally. Since the preferred phrase can vary depending on the sentence context, we conduct ten comparisons for each pair of the 42 hedging phrases, where in each comparison, we randomly sample a sentence in which the phrase occurs. We ask each LLM to identify which phrase conveys greater certainty about the presence of a finding, ensuring that such phrases will eventually receive a higher $\mu$. The main portion of the prompt is shown in Listing 2; the complete prompt additionally includes a system instruction, detailed task description, and three in-context examples (Appendix A.2). To ensure a more neutral comparison, the referenced finding $f_{i}$ within each sentence is masked as <finding>.
@@
With the reference ranking established, we fit each of the 2,066 tentative finding-sentence pairs from the Lunguage dataset into it. Note that we cannot simply use the hedging phrases extracted from the sentence to assign a rank directly, because (i) some phrases are too rare to appear in the reference vocabulary, and (ii) the context in which a phrase occurs can significantly alter its implied certainty. Here, we once again draw inspiration from competitive gaming: when a new player enters a game, TrueSkill quickly estimates their rank by identifying existing players with similar skill levels $\mu$. This is done by selecting opponents with the highest draw probability, iteratively playing those games, and updating the skill level of the new player based on the outcome of each game (Herbrich et al., 2006).
@@ 4.  Implicit uncertainty
Radiology reports often encode diagnostic reasoning implicitly (Gunderman, 2009; Lee et al., 2013; Berlin, 2000; Gärtner et al., 2020), documenting high-level diagnoses while omitting intermediate supporting findings. Although such abstraction enhances efficiency, it introduces ambiguity arising from unmentioned findings. To resolve this, we propose a Pathway Expansion Framework that reconstructs omitted diagnostic evidence by systematically expanding structured findings along predefined diagnostic pathways. This framework yields Lunguage++, an expanded version of Lunguage, integrating explicit and implicit reasoning into a unified structured representation (Figure 4).
@@ 4.1.  Diagnostic Pathway Construction
We formalize radiologists’ implicit diagnostic reasoning into explicit, machine-friendly formats that describe how each diagnosis decomposes into its characteristic radiographic findings. To preserve interpretability and clinical fidelity, we established three principles through expert consensus. (i) Exclusive mutual independence ensures that each pathway differs from others by at least one defining piece of evidence. (ii) Specificity ensures that observations directly contributing to diagnostic differentiation are included, emphasizing findings that are concrete and fine-grained. (iii) High-certainty features retain only clinically consistent and reliable findings to minimize ambiguity and preserve clarity. Pathways were then constructed through a two-stage expert-in-the-loop process grounded in these principles. A radiologist and an AI researcher first defined subfindings for 14 common CXR diagnoses, referencing established radiological interpretation principles (Goodman, 2014; Webb and Higgins, 2011) to identify diagnostic criteria, imaging patterns, and hierarchical relations among findings. These preliminary pathways were subsequently reviewed and refined through consensus by another radiologist and an oncologist.
Pathways are represented as Directed Acyclic Graphs (DAGs) (Pearl, 2009), capturing how higher-level diagnoses require lower-level manifestations within hierarchical reasoning. An example of a DAG corresponding to (part of) the diagnostic pathways for congestive heart failure (CHF), pulmonary edema, pleural effusion, and cardiomegaly is shown in Figure 4. Each node corresponds to an observable finding, and directed edges $(u,v)$ represent diagnostic relations $p_{path}(v|u)$ established through radiologist consensus, indicating that the sub-finding $v$ is typically observed when its parent $u$ is present, with likelihood greater than 0.8. This top-down structure models the cascading sequence of radiographic evidence that underlies diagnostic inference. Distinct pathways are defined for each diagnosis and further refined by imaging view and patient position, since these factors influence the findings one expects to observe. Table 3 reports statistical summaries of the diagnostic pathways, while the full set of pathways is provided in Appendix B.1 and on our Github repository.
@@ 4.2.  Pathway Expansion Framework
Building upon the diagnostic pathways, we design a Pathway Expansion Framework that implements the top-down reasoning process of radiological interpretation to fill in missing sub-findings. Given structured findings from Lunguage (Moon et al., 2026), the framework reconstructs omitted intermediate findings by expanding each diagnosis along its predefined diagnostic pathway, following the recursive procedure detailed in Algorithm 1. It consists of four stages, each of which ensure that reconstructed structures remain clinically faithful and logically consistent.
@@ 1. Finding Deduplication
To ensure that subsequent reasoning operates on a consistent and non-redundant set of findings, overlapping or synonymous mentions referring to the same diagnostic concept at the same location (e.g., “opacity in the basal segment of the lung” and “lung base opacity”) are identified and merged within each report. Pairwise cosine similarity is computed between clinical embeddings (BioLORD (Remy et al., 2023)) of all report findings, where each finding is linearized into a phrase combining its “entity, location, attributes” following the LunguageScore (Moon et al., 2026) formulation. Pairs exceeding a similarity threshold of 0.9 are merged into a single finding, while manually defined blacklist pairs (e.g., left vs. right) are excluded to prevent erroneous merges of semantically distinct findings. This precision-oriented threshold and manual blacklist correction minimize erroneous merges while ensuring that each finding is represented only once before expansion. The detailed blacklist pairs are provided in Appendix B.2.1.